y=4xx+y=5 how do i help my son solve whether it has one solution, no solution or many?
Answer by ewatrrr(24785)

 
Hi,
If you meant:
y = 4x and
x + y = 5
x + 4x = 5 |Substituting the 4x for y
5x = 5
x = 1 and y = 4 (y = 4x = 4·1)
ordered pair (1,4) is the Solution for this system